VOTable Format, Version 1.1 Version 1.05

IVOA Working Draft

2003-11-05

Previous versions:

1.0 (2002-04-15)

Authors:

S. Derriere - derriere@astro.u-strasbg.fr

Introduction

This document is a short example of IVOA reference document in L^{A}TEX. It is designed to produce a PDF file using pdflatex. It should be possible to produce PS output with minor adjustments.

See the source of the template file for details.

Contents

1 Compiling

1.1 Preparing your document

The ivoa latex class is based on the standard article class. To format your document in IVOA style, simply include in the preamble: \documentclass{ivoa}

instead of \documentclass{article}

The standard header is produced with the standard \maketitle command. This command takes a few new options, that are defined in the preamble of the tex file:

• \ivoatype
• \version
• \urlthisversion
• \urllastversion
• \previousversion

See the file template-ivoa.tex for more details.

1.2 Compiling with pdflatex

You can produce PDF directly from the TeX source with pdflatex (or pdftex):

pdflatex myfile.tex

Run it several times for proper cross-references, table of contents, etc. For producing PDF thumbnails, use:

thumbpdf myfile.pdf

This produces a .tpt file. Run pdflatex once more to include them.

1.3 Compiling with L^{A}TEX

You should first make an EPS version of the IVOA logo: convert IVOA.jpg IVOA.eps With latex, you can produce a PS file, and then convert it to PDF.

latex myfile.tex
dvips -o myfile.ps myfile.dvi
ps2pdf myfile.ps myfile.pdf

2 Including figures

2.1 With pdfL^{A}TEX

The best way to include figures is not to specify the file extension:

\includegraphics{my-figure}

pdflatex will search for files with appropriate extensions: PDF, JPEG and PNG can be directly included.

Figure 1: Data processing in astronomy.

PS/EPS files are not supported by pdflatex. However, it is possible to include EPS files with the epstopdf command/package. You must either:

• convert all EPS files into PDF before running pdflatex;

• use the epstopdf package in the preamble, and compile with:

     pdflatex -shell-escape myfile.tex

Note that this will create on-the-fly the .pdf files in your directory ! On next runs of pdflatex, the newly created pdf file will be used if you used \includegraphics{my-figure}. It will be rebuilt each time if

     you specified the extension \includegraphics{my-figure.eps}.

Fig. ?? is an included PDF file.

2.2 With L^{A}TEX

If you really want to take the long road TEX - DVI - PS - PDF, then you cannot include PDF, JPEG or PNG directly.

The IVOA.jpg logo needs to be converted to eps (see section 1.3).

2.3 Examples

Uncomment examples below to test inclusion of various files.

3 Other features

You can include URLs in your document, using the \url{http://www.ivoa.net/} command. It will give direct links in the PDF file if you use the hyperref package in the preamble: http://www.ivoa.net/